Please show work So i can see.

The vapor pressure of pure water at 63.0 C is 175 torr. What is the vapor pressure of an aqueous solution of glucose (a non-volatile nonelectrolyte) at the same temperature if the mol fraction of glucose in the solution is 0.17?
a. 30.0 torr

b. 45.0 torr
c. 226 torr
d. 145 torr
e. We need to know the vapor pressure of pure glucose at that temperature to provide an

The vapor pressure of pure water at 63.0 C is 175 torr. What is the vapor pressure of an aqueous solution of glucose (a non-volatile nonelectrolyte) at the same temperature if the mol fraction of glucose in the solution is 0.17?
a. 30.0 torr

b. 45.0 torr
c. 226 torr
d. 145 torr
e. We need to know the vapor pressure of pure glucose at that temperature to provide an

answer.

answer.
Answer #1

Given,

mole fraction of Glucose,

X(Glucose) = 0.17

So,

X(solvent) = 1 - X(Glucose)

= 1 - 0.17

= 0.83

According to Raoult’s law:

P = Po*X(solvent)

p = 175*0.83

p = 145 torr

Answer: d